B.I.D.O.N Rue des Mots 80886 Sassonne-le-Creux

T E L E F A X

TELEPHONE: 987.64.20 TELEFAX: 987.75.31

To: Geneva University To whom it may concern

Telefax: +41-22-320 29 27

From: Dr D. Mégevand TEXnician

Total pages: 1

If improperly transmitted, please call the telefax operator

Sassonne-le-Creux, January 7, 2002

Subject: L^{A}TEX fax style

Dear collegues

This is just an example of my new telefax style in english. See the french example for the documentation.

Looking forward to hearing from you soon and thanking you in advance.

Sincerely yours,

Denis Mégevand

E-Mail:  megevand@obs.unige.ch                    megevand at cgeuge54